OpenStudy (anonymous):

A number decreased by negative three is equal to -40. Write an equation and find the number.

OpenStudy (anonymous):

x+3=-40

OpenStudy (anonymous):

Let the number be x. x - 3 = -40 x = -40 + 3 = -37

OpenStudy (anonymous):

x= -43

OpenStudy (anonymous):

decreased by negative three means x -(-3)

OpenStudy (anonymous):

oh yes I thought decreased as increased.

OpenStudy (anonymous):

oh.............

OpenStudy (anonymous):

thanks for correcting my mistake.

OpenStudy (anonymous):

welcomes......

OpenStudy (anonymous):

The answer is -43.

OpenStudy (anonymous):

hahahahaha

OpenStudy (anonymous):

let the no. be x x-(-3)=-40 x=-40+(-3) x=-43 simple..

OpenStudy (anonymous):

let x be the number therefore x-3=-40 then by adding 3 to both side we havw x-3+3=-40+3 x=-37